# Training Plug-n-Play Knowledge Modules with Deep Context Distillation

### v1

@@ 4 Related Work
Knowledge Distillation typically transfers logits or intermediate representations from a teacher model to a student (Xu et al., 2024). While effective at compressing large models, these approaches usually differ from KMs in that they do not explicitly store domain knowledge for repeated use in specific tasks. Context distillation (Snell et al., 2023) aims to “absorb” a context into a model in a way that preserves its impact on generation. Early work focused on toxicity and safety contexts (Askell et al., 2021) and for internalizing reasoning traces (Snell et al., 2023). Recently, Shin et al. (2024) propose Generative Context Distillation, augmenting the distillation objective with a generative loss to match model outputs more comprehensively.
@@
Knowledge Injection with Modules Several works have similar goals of KMs. Xiao et al. (2023) introduce “plug-and-play” document modules where an MLP transforms encoded document tokens into soft prompts used across attention layers. Zhang et al. (2023) similarly train knowledge modules in a task-agnostic or task-specific manner, caching the processed document representation for efficient reuse. Amortization-based approaches train a “hypernetwork” to generate lightweight adaptations (e.g., low-rank shifts or prefix-tuning parameters) from a given context. Chen et al. (2024) learn to project context tokens into low-rank parameter shifts in a self-supervised manner, and Tack et al. (2024) encode documents into prefix-tuning weights using a T5-based hypernetwork. These methods train the hypernet on multi-task data, so at inference time it can produce task-specific or document-specific modules in a single forward pass. KMs as described in this paper are trained with gradient descent on single documents independent of any multi-task training. This increases per-document costs but reduces the risk of domain shift. Future work might be devoted to efficient learning of KMs.
@@ 5 Conclusion
In this paper, we proposed an alternative approach to the de-facto next-word prediction loss for knowledge extraction of documents; we showed that using Deep Context Distillation, with synthetic data, yields consistent improvements across several question-answering benchmarks. Furthermore, we showed that our approach pairs nicely with RAG, offering practitioners multiple options to maximize performance, depending on their inference budget.

### v2

@@ 4 Related Work
Knowledge Distillation typically transfers logits or intermediate representations from a teacher model to a student (Xu et al., 2024). While effective at compressing large models, these approaches usually differ from KMs in that they do not explicitly store domain knowledge for repeated use in specific tasks. Context distillation (Snell et al., 2023) aims to “absorb” a context into a model in a way that preserves its impact on generation. Early work focused on toxicity and safety contexts (Askell et al., 2021) and for internalizing reasoning traces (Snell et al., 2023). Recently, Shin et al. (2024) propose Generative Context Distillation, augmenting the distillation objective with a generative loss to match model outputs more comprehensively. Choi et al. (2023) use generative distillation to internalize long prompts, first training a generator on data sampled from the prompt-conditioned LLM, and subsequently using the generator’s samples for the distillation step. In contrast to these approaches, the work presented in this paper goes beyond prompt internalization, and investigates distillation as a mechanism for knowledge acquisition. Recently, Qi et al. (2025) show that a distillation objective can be used for localized knowledge editing given (entity, relation, object) triplets. Finally, concurrent to our work, Kujanpää et al. (2024) also explore the synergy between knowledge distillation and synthetic data generation. Our work still differs by the modularity or our solution, which consists in learning a KM for each document and demonstrate transferability of the solution across unseen KMs. Finally, our work goes beyond self-distillation, as we show that using synthetic data from a stronger (but off-policy) model yields additional gains. See Tab. 3 for a summary.
Adapter Loss Synthetic Module
Granularity Function Data Composition
Xiao et al. (2023) Document LM, Masked LM No Yes
Zhang et al. (2023a) Corpus Masked LM No Yes
Kujanpää et al. (2024)a Corpus Distillation Yes No
Knowledge Modules Document Distillation Yes Yes
  • a
    Concurrent work
Table 3: Comparison of Knowledge Injection Methods with respect to our proposed approach, Knowledge Modules.
@@
Knowledge Injection with Modules Several works have similar goals of KMs. Xiao et al. (2023) introduce “plug-and-play” document modules where an MLP transforms encoded document tokens into soft prompts used across attention layers. Unlike the work in Xiao et al. (2023), we require that the inference compute is fixed and does not scale with the length of the document $D$Zhang et al. (2023a) similarly train knowledge modules in a task-agnostic or task-specific manner, caching the processed document representation for efficient reuse. Amortization-based approaches train a “hypernetwork” to generate lightweight adaptations (e.g., low-rank shifts or prefix-tuning parameters) from a given context. Chen et al. (2024) learn to project context tokens into low-rank parameter shifts in a self-supervised manner, and Tack et al. (2024) encode documents into prefix-tuning weights using a T5-based hypernetwork. These methods train the hypernet on multi-task data, so at inference time it can produce task-specific or document-specific modules in a single forward pass. KMs as described in this paper are trained with gradient descent on single documents independent of any multi-task training. This increases per-document costs but reduces the risk of domain shift. Future work might be devoted to efficient learning of KMs.
KV Cache Compression Our work can be seen as an instantiation of context compression (Nawrot et al., 2024). Knowledge Modules operate in the query-agnostic setting (Zhang et al., 2023b; Devoto et al., 2024a), where compression is performed without knowledge of the downstream task. Several approaches can also train the context compressor, typically with cross-entropy, or a reconstruction objective (Ge et al., 2024; Qin et al., 2024) More closely related to ours, Chari et al. (2024) leverage a distillation objective to perform prompt compression into a soft-prompt-like adapter.
@@ 5 Conclusion
In this work, we introduced a novel approach to encode document-specific knowledge into lightweight, plug-and-play Knowledge Modules (KMs) using Deep Context Distillation (DCD). We showed that DCD, especially paired with synthetic data, outperforms traditional next-token prediction and pre-instruction tuning across several question-answering benchmarks. Our results highlight DCD’s consistent performance gains, its synergy with retrieval-augmented generation (RAG), and the modularity that enables fine-grained control over knowledge integration.
The current investigation focused on the setting where the document pertaining to the question is known. For future work, we are excited by the potential combination of KMs with zero-shot routing methods (Ostapenko et al., 2024; Muqeeth et al., 2024) to deploy our approach in document-agnostic settings, where the system can combine information from multiple KMs.
Appendix A Synthetic Data Generation
Below we provide additional information on the generated data used for our experiments. We split documents into non-overlapping chunks of 2048 tokens, using RecursiveCharacterTextSplitter from the langchain_text_splitters package. The prompts are provided below.
A.1 Summary Generations
We generate a maximum of 16 summaries for each chunk. For our main experiments, we use a default of 8 summaries per chunk, unless stated otherwise.
Summary Generation Prompt Summarize the following text in around {int(len(text) / 4)} words without omitting any important details. The summary should be grammatically correct and summarize all the different sections in the text
********** Text **********
{text}
A.2 Q/A Generation
Q/A Generation Prompt Create num_questions questions that can be answerable from the following text, along with their answers. Strive to generate challenging questions that require aggregating information across the provided text. Focus on different sections of the text to increase diversity of the generated questions. Format your answer as follows:
<question id=1 >QUESTION 1 HERE <question >
<answer id=1 >ANSWER 1 HERE <answer >
<question id=2 >QUESTION 2 HERE <question >
<answer id=2 >ANSWER 2 HERE <answer >
********** Text **********
{text}
A.3 Entigraph Generation
For the Entigraph data, we use the publicly available Entigraph generated dataset on the QuALITY dataset available on HuggingFace under zitongyang/entigraph-quality-corpus. The dataset provised 1 (large) synthetically generated document per document in the original dataset. In order to obtain smaller chunks of synthetic data, we split according to <|entigraphseptoekn|>, which gives us chunks of 7̃00 tokes on average describing a specific relationship between two entities. As we do not know where the information is contained in the original document, we provide the full document in context during training.
Appendix B Ablations
The following ablations are done on a 140 document subset of NQA (100 train, 20 dev, 20 test).
Method Rouge-L (no KE) Rouge-L (KE)
KM (Summary DCD) 23.1 32.3
KM (Summary DCD, hidden loss only) 18.8 29.5
KM (Summary DCD, logit loss only) 21.7 32.3
Table 4: Ablation on the each distillation component using Phi-3 as the backbone.
Method Rouge-L (no KE) Rouge-L (KE)
KM (2 summaries) 20.4 31.0
KM (4 summaries) 21.8 31.5
KM (8 summaries) 23.2 31.7
KM (16 summaries) 24.4 32.3
KM (Q/A pairs) 22.1 31.7
KM (summaries & Q/A pairs) 26.3 32.7
Table 5: Ablation on the data used for DCD training